OpenStudy (anonymous):

What is the equivalent resistance in a parallel circuit that has the following three resistors: 6.8 ohms, 9.0 ohms, and 15.0 ohms? A. 10.3 ohms B. 30.8 ohms C. 3.08 ohms D. 0.80 ohms E. 12.8 ohms

OpenStudy (binary3i):

\[(1/Req) = (1/R1)+(1/R2)+(1/R3)\] use this

OpenStudy (anonymous):

u divide each one by 1?

OpenStudy (anonymous):

and add them together

OpenStudy (anonymous):

whats the req?

OpenStudy (souvik):

equivalent Resistance is req

OpenStudy (souvik):

1/R=1/6.8+1/9.0+1/15 find R

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

3.8

OpenStudy (anonymous):

3.08

OpenStudy (anonymous):

u just round it right

OpenStudy (anonymous):

3.075

OpenStudy (souvik):

3.08...

OpenStudy (anonymous):

thanks:)